OpenStudy (anonymous):

Area of a circle is 907.9 square inches. Find the radius.

OpenStudy (anonymous):

\[\large A=\pi r^2\] A=Area of a circle r=radius

OpenStudy (johnweldon1993):

Area of circle \[\pi r^2\] You have Area so \[\pi r^2 = 907.9\] Solve for 'r'

OpenStudy (anonymous):

Got it. (: Thank you '

OpenStudy (johnweldon1993):

no problem!